If 2.6m3 of a gas initially at STP is compressed to 1.2m3 and its temperature is raised to 40?C, what is its final pressure? (answer in atm)
Answer #1

Note that

P1 V1 / T1 = P2 V2 / T2

As

P1 = 1 atm
V1 = 2.6 m^3
V2 = 1.2 m^3
T1 = 273.15 K
T2 = 313.15 K

Then

P2 = 2.48 atm   [ANSWER]
Answer #2

P1=1 atm,

V1= 2.6 m3

T1= 273K

P2=?

V2= 1.2 m3

T2= 313K

P1V1/T1= P2V2/T2

P2 = P1V1T2/V2T1

= (1X2.6 X313)/(1.2X273)=2.484 atm